[Halla_running] [New Logentry] RC Daily Summary
adhidevi at jlab.org
adhidevi at jlab.org
Sat Aug 16 09:50:02 EDT 2025
Logentry Text:
--
We are in steady production. The beam is good, and the experiment is running well, except for a few DAQ crashes (https://logbooks.jlab.org/entry/4443532) and a few GEM trips (L15M0 and L15M1). The SBS magnet corrector issue was resolved relatively quickly yesterday morning (https://logbooks.jlab.org/entry/4443076). The plan for the weekend is to continue production with a 15 μA beam.
Charge collected since yesterday's Day to today's owl:
* Total Accumulated Charge at Kin3 so far: 84.06 C (Goal: 100 C)
* Total Accumulated Charge at Kin3 since yesterday : 0.93 C
Shift Summary:
* Yesterday's Day: https://logbooks.jlab.org/entry/4443015
* Yesterday's Swing: https://logbooks.jlab.org/entry/4443403
* Today's Owl: https://logbooks.jlab.org/entry/4443555
There is no RC Meeting for today. Have a good weekend.
---
This is a plain text email for clients that cannot display HTML. The full logentry can be found online at https://logbooks.jlab.org/entry/4443737
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/halla_running/attachments/20250816/67ed391e/attachment.htm>
More information about the Halla_running
mailing list